1/3 multiplied by 452.16
Found 2 solutions by themasterofcircuits, josgarithmetic:
Answer by themasterofcircuits(27)

1/3*452.16=0.33*452.16= [149.21]

Answer by josgarithmetic(39617)

Long division of 452.16 divided by 3.



150.72